The rotational constant and centrifugal correction factors for HBr are 8.473 cm-1 and 0.0003 cm-1 , respectively:

(a) What are the energies of the 3 lowest
Answer #1

lowest possible value of J is 0

given

rotational constant B=8.473 cm-1)

centrifugal constant D=0.0003 cm-1

Energy E0= BJ(J+1) -DJ2(J+1 )2 cm-1

for  J=0

E0= 8.473 *0 (0+1) - 0.0003*0(0+1 )2 cm-1

E0 = 0

next level

J=1

Energy E1= BJ(J+1) -DJ2(J+1 )2 cm-1

J=1

E1= 8.473 *1(1+1) - 0.0003*12(1+1 )2 cm-1

E1= 8.473 *2 - 0.0003*4 cm-1

E1 = 8.473 *2 - 0.0003*4 ​​​​​​​ cm-1

E2 = 16.946-1.2*10-3​​​​​​​ cm-1

E1 = 16.9448​​​​​​​ cm-1

next level

J=2

Energy E1= BJ(J+1) -DJ2(J+1 )2 cm-1

E2= 8.473 *2(2+1) - 0.0003*22(2+1 )2 cm-1

E2= 8.473 *6 - 0.0003*4*9 cm-1

E2 = 8.473 *6 - 0.0003*36 ​​​​​​​ cm-1

E2 = 50.838 - 0.0108 ​​​​​​​ cm-1

E2= 50.8272​​​​​​​ cm-1

the energies of the 3 lowest level of HBr are 0,16.9448​​​​​​​ cm-1, 50.8272​​​​​​​ cm-1